OpenStudy (anonymous):

A game is played where 26 tiles are in a bag. Each tile has a different letter of the alphabet on it. If you draw four letters, what is the probability of the letters being M-A-T-H in order, if there is no replacement?

OpenStudy (anonymous):

1/26*1/25*1/24*1/23=1/358800=.000002787